A charged particle moves in a plane perpendicular to a magnetic field B. If it starts with some additional speed v parallel to B, the trajectory of the charged particle will be
1
Expert's answer
2019-05-14T10:10:04-0400

The trajectory is a spiral, its axis is parallel to the magnetic field. The step of the spiral is


"h = v \frac{2\pi}{\omega} = v \frac{2\pi m}{eB}"

in nonrelativistic mechanics and


"h = v \frac{2\pi E}{eBc^2}"

in relativistic mechanics.